OpenStudy (anonymous):

h(x)=2x-5/3 find it inverse function

OpenStudy (unklerhaukus):

is that \[h(x)=2x-\frac53\] or\[h(x)=\frac{2x-5}3\]

OpenStudy (unklerhaukus):

in either case solve the equation for x, Then switch x for h and h for x, done.

OpenStudy (anonymous):

\[h(x)=\frac{ 2x-5 }{ 3 }\]

OpenStudy (unklerhaukus):

so can you solve it for x=? (multiply both sides by 3, add 5 to both sides, then divide both sides by 2)

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

do i have to change the h(x) into y?

OpenStudy (unklerhaukus):

you can if you like

OpenStudy (unklerhaukus):

if you were going to do that you should make it the first step,

OpenStudy (anonymous):

\[h(x)=\frac{ 3x+5 }{ 2 }\] in this the answer?

OpenStudy (anonymous):

is this the answer? sorry wrong grammar

OpenStudy (unklerhaukus):

yeah that's the inverse

OpenStudy (unklerhaukus):

good work

OpenStudy (anonymous):

thanks !